OpenStudy (anonymous):

What is the middle term of the product of (x + 5)(x + 2)?

OpenStudy (primeralph):

What do you think?

OpenStudy (primeralph):

Multiply out.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

FOIL; First Outer Middle Last

hero (hero):

(x + 5)(x + 2) = x(x + 2) + 5(x + 2) = x^2 + 2x + 5x + 10 = x^2 + 7x + 10 The middle term is 7x

OpenStudy (primeralph):

@Hero What's up with the answers?

hero (hero):

She asked "how". Maybe she never seen it multiplied before. Seeing the process helps.

hero (hero):

I showed her the steps for how to get from the problem to the solution.

OpenStudy (primeralph):

Well, you answered another one of her questions earlier the same way but never checked if she actually understood it. That might be better for her.

hero (hero):

If she has any questions concerning how to arrive at the solution or if she is confused about anything it is up to her to ask so that I can clarify such things.

hero (hero):

I'm not her tutor or teacher. I'm just here to assist.

OpenStudy (primeralph):

As you like it.